OpenStudy (volleyballlover55):

Who can help me with Pre-Algebra!! :) Make sure you answer in your OWN words please!

OpenStudy (volleyballlover55):

Describe the difference between rational and irrational numbers and give an example of each.

OpenStudy (volleyballlover55):

@Mehek14 @rebeccaxhawaii @anthonyym

OpenStudy (anthonyym):

A rational number can be expressed as a fraction while an irrational number cannot.

OpenStudy (volleyballlover55):

ok and can you give an example for each

OpenStudy (savanna_18):

a rational number can be written as a ratio, an irrational number can be written as a decimal but not as a fraction.

OpenStudy (volleyballlover55):

ok and give an example for each please!!

OpenStudy (anthonyym):

Example of irrational: pi Example of rational: 5/3

OpenStudy (savanna_18):

8 is rational because it can be put in a fraction 1/8 pie (3.14) would be irrational because it goes on and are non-repeating digits to the right of the decimal point. 3.141592..... and so on

OpenStudy (volleyballlover55):

8/1 =8 not 3.14

OpenStudy (volleyballlover55):

i m sorry i mean .. 1/8=0.125

OpenStudy (savanna_18):

what i was trying to say was 3.14 is irrational and 8 can be put in a fraction as 8/1so that would make it rational.

OpenStudy (volleyballlover55):

oh ok.. awesome! next question!

OpenStudy (savanna_18):

okay what is it?

OpenStudy (volleyballlover55):

Approximate the square root of 110.

OpenStudy (volleyballlover55):

I'm going to leave but ill be back in about 1 hour... :)

OpenStudy (savanna_18):

that's it just 110 you can't do more to it. (:

OpenStudy (volleyballlover55):

nvm im not going to leave.

OpenStudy (anthonyym):

To approximate the square root of 110, you can type it into the calculator and round it to a number of digits.

OpenStudy (volleyballlover55):

wait so tell me the ANSWER.. Can it be approximate or no?

OpenStudy (volleyballlover55):

??

OpenStudy (volleyballlover55):

@savanna_18

OpenStudy (savanna_18):

i just got confused ask again please?

OpenStudy (volleyballlover55):

Approximate the square root of 110.

OpenStudy (savanna_18):

So you want the approximate answer of the squared root of 110?

OpenStudy (volleyballlover55):

yes :)

OpenStudy (volleyballlover55):

@savanna_18 ?

OpenStudy (savanna_18):

okay sorry, i believe you can't do anything more this answer i believe the answer would be left at 110 from my understanding. i'm not to sure about this one.

OpenStudy (volleyballlover55):

i know it can't be simplified anymore but i want to know why.. :) can you explain it to me :)

OpenStudy (volleyballlover55):

@savanna_18 ????

OpenStudy (savanna_18):

the square root of 110 is in simplest form since 11= 11*5*2 (none are square) consider 50=5*5*2=5^2*2 so sqrt (50)=5* sqrt(2) Does this help?

OpenStudy (volleyballlover55):

yes :)

OpenStudy (savanna_18):

okay great (:

OpenStudy (volleyballlover55):

how do you find the approximation of a square root?

OpenStudy (volleyballlover55):

like with any number .. ? like if my teacher asks me .. Approximate the square root of any number .. how do i find it?

OpenStudy (savanna_18):

Find each approximate answer for each square root. You round your decimal answer to the nearest tenth.

OpenStudy (volleyballlover55):

ok next!

OpenStudy (savanna_18):

okay

OpenStudy (volleyballlover55):

What is the cubed root of 27?

OpenStudy (savanna_18):

3

OpenStudy (volleyballlover55):

ok and can you tell me the steps you used to get that answer..

OpenStudy (savanna_18):

3^3= 27

OpenStudy (anthonyym):

I think there's a difference between the approximation of a square root and simplifying the square root. For example, sqrt8 approximates to 2.83, but simplifies to 2*sqrt2

OpenStudy (volleyballlover55):

right.... i need to approximate the square root of 110. @anthonyym

OpenStudy (anthonyym):

So your directions say "Find each approximate answer for each square root. You round your decimal answer to the nearest tenth." Notice it says to round the decimal answer to the nearest tenth. To get a decimal answer, simply type \[\sqrt{110}\] into the calcalator, in which you would get 10.48808848, and rounding to the nearest tenth, it would be 10.5

OpenStudy (volleyballlover55):

well the question is Approximate the square root of 110. @anthonyym

OpenStudy (volleyballlover55):

is it the same answer.. ?

OpenStudy (anthonyym):

No, the approximation 10.5 is not the same as writing\[\sqrt{110}\], because an approximation is an estimate.

OpenStudy (volleyballlover55):

so the answer is the approzimation of the square root of 110 is 10.5?

OpenStudy (volleyballlover55):

@anthonyym

OpenStudy (anthonyym):

Yes

OpenStudy (volleyballlover55):

ok :)

OpenStudy (volleyballlover55):

@anthonyym